JuliMa:

A current is passed through a Ga(NO3)3 solution for 1.5 hours, and after this time period the mass of metal produced was 6.3 grams. What is the current, in amperes, that is required to produce such an amount of gallium?

supie:

Choices would help...

supie:

idk much about this but I can try.... \[Ga^3+ + 3e \ \ ➡\ Ga\] A unit electric charge (faraday) and so \[1\ faraday= 96500C\]

supie:

We have to do the value of 1 faraday multiplied by three \[3\ faraday = 3* 96500C = 289500C\]

supie:

According to \[Ga^3++3e \ ➡ Ga\]\[70g/Ga\] was deposited by 289500C meaning that for this situation \[6.3g/Ga\] will now be deposited \[(6.3 x 289500)/6.3 = 26055C\]

supie:

Now we calculate the `current flow` \[t = 1.5\ hours = 1.5 * 3600 = 5400\ seconds\]

supie:

\[Q = 26055C\]\[Q = It\]\[I\ =\ Q/t\]\[I\ =\ 26055/5400\] \[I \ =\ ?\]